# Grappa - A Machine Learned Molecular Mechanics Force Field

### v1

@@ Abstract
Simulating large molecular systems over long timescales requires force fields that are both accurate and efficient. In recent years, E(3) equivariant neural networks have lifted the tension between computational efficiency and accuracy of force fields, but they are still several orders of magnitude more expensive than classical molecular mechanics (MM) force fields. Here, we propose a novel machine learning architecture to predict MM parameters from the molecular graph, employing a graph attentional neural network and a transformer with symmetry-preserving positional encoding. The resulting force field, Grappa, outperforms established and other machine-learned MM force fields in terms of accuracy at the same computational efficiency and can be used in existing Molecular Dynamics (MD) engines like GROMACS and OpenMM. It predicts energies and forces of small molecules, peptides, RNA and — showcasing its extensibility to uncharted regions of chemical space — radicals at state-of-the-art MM accuracy. We demonstrate Grappa’s transferability to macromolecules in MD simulations, during which large protein are kept stable and small proteins can fold. Our force field sets the stage for biomolecular simulations close to chemical accuracy, but with the same computational cost as established protein force fields.
@@ 1 Introduction
Machine learned force fields are currently reshaping the field of computational chemistry and Molecular Dynamics (MD) simulations. E(3)-equivariant neural networks [1, 2, 3, 4, 5] are capable of predicting energies and forces of small molecules to great accuracy with lower computational cost than quantum mechanical (QM) methods. However, these models are several orders of magnitude more expensive than Molecular Mechanics (MM) force fields, which employ a simple physics-inspired functional form to parametrize the potential energy surface of a molecular system, hence trading off accuracy in favor of efficiency. For MD simulations of large systems such as proteins and nucleic acids, MM force fields are well established and widely used. Traditional MM force fields rely on lookup tables with a finite set of atom categories that describe local chemical environments for parameterization. Recently, the Espaloma approach [6] has demonstrated that machine learning can be used to increase the accuracy of MM force fields by learning to assign parameters based on a graph representation of the molecule with hand-crafted chemical features such as orbital hybridization states.
@@
In this work, we propose a novel machine learning architecture, Grappa (Graph Attentional Protein Parametrization), to learn MM parameters directly from the molecular graph, improving accuracy on a broad range of chemical space and eliminating the need for hand-crafted features. Grappa employs a graph attentional neural network to predict local chemical environments followed by a transformer [7] with symmetry-preserving positional encoding. Since MM parameters only have to be predicted once per molecule, Grappa can be incorporated into highly optimized MM engines such as GROMACS [8] and OpenMM [9]. This allows energy and force evaluations with the same computational cost as traditional force fields, at state-of-the-art MM accuracy.
@@
We show that Grappa outperforms traditional MM force fields and Espaloma on the Espaloma dataset [10], which contains over 14,000 molecules and more than 1 million states, covering small molecules, peptides and RNA. Since Grappa uses no hand-crafted chemical features, it can be extended to uncharted regions of chemical space, which we demonstrate on the example of peptide radicals. Grappa is transferable to macromolecules such as proteins, which it keeps stable during MD over nanosecond timescales. Starting from an unfolded initial state, MD simulations of small proteins parametrized by Grappa recover experimentally determined folding structures of small proteins, suggesting that Grappa captures the physics underlying protein folding.
@@ 1.1 Molecular Mechanics
Given a molecular graph, MM expresses the potential energy of a system as a sum of contributions from different interactions. Bonded interactions are described by functions of E(3)-invariant internal coordinates such as the lengths $r_{ij}$ of bonds between two atoms, angles $\theta_{ijk}$ between three consecutive atoms and dihedrals $\phi_{ijkl}$ of two planes spanned by four atoms. For the dihedrals, one considers interactions between four atoms that are either consecutively bonded (torsions) or where three atoms are bonded to a central atom (impropers), which do not reflect an independent degree of freedom but are used to maintain planarity of certain chemical groups. One commonly uses harmonic potentials for bond stretching and angle bending and a periodic function for the dihedral potential. The potential energy of all interactions along bonds then is given by
@@
For the dihedral potential in Grappa, we choose a Fourier series with the constraint that the dihedral potential is extremal at and symmetric around zero, which eliminates the need for sine terms. Additionally, atom pairs that are not included in such N-body bonded interaction terms contribute to the potential energy through pairwise nonbonded interaction, typically described by Lennard-Jones and Coulomb potentials.
@@ 2 Grappa
Inspired by the atom typing with hand-crafted rules in traditional MM force fields and in analogy to [6], Grappa first predicts $d$-dimensional atom embeddings,
@@
Since the mapping from molecular graph to energy is differentiable with respect to the model parameters and spatial positions, it can be optimized on predicting QM energies and forces end-to-end. Notably, the machine learning model does not depend on the spatial conformation of the molecule, thus it has to be evaluated only once per molecule and the computational cost of each subsequent energy evaluation is given by the MM energy function.
Grappa currently only predicts bonded MM parameters since the highly nonlinear energy function for nonbonded interactions impedes optimization. The nonbonded parameters are taken from established MM force fields that can reproduce solute interactions and melting points, which we expect to be strongly dependent on nonbonded interactions.
@@ 2.1 Permutation symmetries in MM
For the mapping from atom embeddings $\bm{\nu}$ to MM parameters, we postulate certain permutation symmetries that our model should respect. To derive these symmetries, we consider the energy function of MM as a decomposition into contributions from subgraphs of the featurized molecular graph that correspond to bonds, angles, torsions and improper dihedrals. We demand invariance of the energy contribution under node permutations that induce isomorphisms of the respective subgraph. For bonds, angles and torsions, these permutations leave the respective spatial coordinate invariant, thus we can simply achieve invariance of the energy contribution by demanding invariance of the MM parameters,
@@ 2.2 The Grappa architecture
To predict atom embeddings from the molecular graph, Grappa employs a graph attentional neural network inspired by the transformer architecture [7]. Multi-head dot-product attention on graph edges [11] is followed by nodewise feed-forward layers. We use residual connections [12] and layer normalization [13], which has been demonstrated to enhance the expressivity of the attention layer [14].
@@
defining a symmetry pooling operation with $\mathcal{P}$-invariant output. We call this combination of permutation invariant positional encoding with permutation equivariant layers and symmetric pooling the symmetric transformer, which is illustrated in Figure 2. The symmetric transformer can be generalized to permutation symmetries of arbitrary subgraphs by using the eigenvectors of the graph Laplacian [15] as positional encoding.
Finally, we map to the range of physically sensible parameters, e.g. $(0,\infty)$ for bond and angle force constants or $(0,\pi)$ for equilibrium angles $\theta^{(0)}$. To this end, we use scaled and shifted versions of ELU and the sigmoid function as described in A.2.3. While one could also use the exponential for mapping to $(0,\infty)$, ELU’s linear behaviour towards large inputs is favorable for producing stable gradients during optimization. With the scaling we ensure that a normally distributed output of the neural network is mapped to a distribution with mean and standard deviation that is suitable for the respective MM parameter, which can be seen as a normalization technique [13]. For predicting dihedral parameters $\xi_{ijkl}$, we use a sigmoid gate, which allows the model to supress dihedral modes that are not needed.
As input feature, we use one-hot encodings of the atomic number, the number of neighbors of the respective node and membership in loops of length 3 to 8, which can be directly calculated from the molecular graph. The assignment of nonbonded MM parameters is done using a traditional force field of choice. Since the optimal bonded parameters predicted by Grappa may depend on the scheme by which nonbonded parameters are assigned, we pass a sinusodial encoding [7] of the partial charge of each atom as input feature to Grappa, which also allows to encode the total charge of a molecule without breaking graph symmetries as one would potentially do by using the formal charge instead. For the extension to proteins in section 3.3, we train Grappa to be compatible with two different schemes of nonbonded parameter assignment, which we encode as additional input feature. These schemes are a) what we call the AM1-BCC scheme, with Lennard-Jones parameters from the OpenFF-2.0.0 force field and partial charges from the semi-empirical method AM1-BCC [16] and b) what we call the ff99SB scheme, with Lennard-Jones parameters and partial charges from the established Amber ff99SB-ILDN [17] protein force field.
@@ 2.3 Training
We train Grappa to minimize the mean squared error (MSE) between QM energies $E_{\text{QM}}$ and forces $-\nabla_{\mathbf{x}}E_{\text{QM}}$, whose contribution we weight by the hyperparameter $\lambda_{F}$, and its prediction, $E$ and $\nabla_{\mathbf{x}}E$ respectively, which includes the non-learnable nonbonded contribution. At the start of training, we also include the deviation of predicted bond, angle and torsion MM parameters to those of a given traditional force field, weighted by $\lambda_{\text{trad}}$ and, for regularization, the $L2$ norm of dihedral MM parameters $\xi_{ijkl}^{(\text{dih})}$, weighted by $\lambda_{\text{dih}}$ as in [10]. That is, we use the loss function
@@
Since MM can only predict energy differences of states, not formation energies, we subtract the mean of target and predicted energies for each molecule. As shown previously for other machine learned force fields [18, 5], we find training on forces $-\nabla_{r}E$ in addition to energies important also in our setting, for two reasons. First, the energy is a global, pooled quantity and thus less expressive than the forces, which are local and by a factor of $3N$ more numerous. Second, as shown in [19], learning derivatives of the target with respect to the input can lead to improved generalization and data efficiency, effectively smoothing the learned potential energy surface.
The motivation for the different terms in the loss function, details on the training procedure and hyperparameters for all models mentioned in Sections 3 and 4 can be found in A.3.
@@ 3 Results
3.1 Benchmark results
To demonstrate that Grappa is the current state of the art MM force field, we train and evaluate a Grappa model on the dataset reported in the follow-up paper [10] to Espaloma, which contains 17,427 unique molecules and over one million conformations. Our training and test partition is identical with the one from Espaloma, where the molecules were divided into 80% training, 10% validation and 10% test set, based on isomeric SMILES strings. The dataset covers small molecules, peptides and RNA with states sampled from the Boltzmann distribution at 300 K and 500 K, from optimization trajectories and from torsion scans. The nonbonded contribution is calculated using the OpenFF-2.0.0 force field [20] and partial charges from the AM1-BCC method, as in Espaloma.
For all types of molecules, Grappa outperforms established MM force fields and Espaloma in terms of energy and force accuracy on Boltzmann-sampled states as shown in Table 1 and Figure 3a. While the Boltzmann samples are the more relevant benchmarking data for using Grappa in MD simulations, we also compared performance on torsion scans and optimization trajectories (Table A4). There, we find Grappa and Espaloma to be competitive, while Grappa is more accurate for forces and less accurate for energies. This is also the case if Grappa is only trained on a fraction of Espaloma’s training set (Figure A9), indicating high data efficiency.
@@ 3.1 Benchmark results
Table 1: Accuracy of Grappa, Espaloma 0.3 [10] and established MM force fields on test molecules of the Espaloma dataset. We only display the states that are sampled from the Boltzmann distribution, optimization trajectories, torsion scans and uncertainties are shown in Table A4. We report the RMSE of molwise-centered energies in kcal/mol and the componentwise RMSE of forces in kcal/mol/Å. Gaff-2.11 [21] is a general-purpose force field, ff14SB [22] is an established protein force field and RNA.OL3 [23] is specialized to RNA. The datasets SPICE-Pubchem and SPICE-DES-Monomers are subsets of the SPICE dataset [24] and contain small molecules, the other datasets contain peptides and RNA. As in Espaloma, we bootstrap the set of test molecules 1000 times and report the mean of the RMSEs; their standard deviation is given in Table A4.
3.2 Latent space
@@
Figure 3: (a) Comparison of energy and force predictions of Grappa (model from Table 1) and established force fields for 10,000 randomly chosen conformations of test molecules from Espaloma’s SPICE-Pubchem, SPICE-Dipeptide and RNA-Trinucleotide dataset. (b) The two first principal components $u_{1}$ and $u_{2}$ of predicted atom embeddings from the Espaloma test dataset can be related to a combination of the main group and period in the periodic table of elements. Lines of constant main group or period are represented by approximate diagonals in latent space.
To generalize across the broad range of chemical space covered by the training set, our model has to learn highly informative atom embeddings that capture the local environment in the molecular graph. For interpreting this latent representation, we apply a two-dimensional principal component analysis on a set of predicted atom embeddings (3) of test molecules (Figure 3b). It turns out that the two principal components can be related to the periodic table of elements, which is thus learned implicitly by Grappa.
3.3 Extension to proteins

### v2

@@ Abstract
Simulating large molecular systems over long timescales requires force fields that are both accurate and efficient. In recent years, E(3) equivariant neural networks have lifted the tension between computational efficiency and accuracy of force fields, but they are still several orders of magnitude more expensive than established molecular mechanics (MM) force fields. Here, we propose Grappa, a machine learning framework to predict MM parameters from the molecular graph, employing a graph attentional neural network and a transformer with symmetry-preserving positional encoding. The resulting Grappa force field outperformstabulated and machine-learned MM force fields in terms of accuracy at the same computational efficiency and can be used in existing Molecular Dynamics (MD) engines like GROMACS and OpenMM. It predicts energies and forces of small molecules, peptides, RNA and — showcasing its extensibility to uncharted regions of chemical space — radicals at state-of-the-art MM accuracy. We demonstrate Grappa’s transferability to macromolecules in MD simulations from a small fast folding protein up to a whole virus particle. Our force field sets the stage for biomolecular simulations closer to chemical accuracy, but with the same computational cost as established protein force fields.
@@ 1 Introduction
In recent years, advances in geometric deep learning have led to the development of highly accurate machine learned force fields, reshaping the field of computational chemistry and Molecular Dynamics (MD) simulations. E(3) equivariant neural networks (Musaelian et al., 2022; Batatia et al., 2023; Unke et al., 2022; Schütt et al., 2021) are capable of predicting energies and forces of small molecules to great accuracy with lower computational cost than quantum mechanical (QM) methods. However, these models are several orders of magnitude more expensive than Molecular Mechanics (MM) force fields, which employ a simple physics-inspired functional form to parametrize the potential energy surface of a molecular system, hence trading off accuracy in favor of efficiency. For MD simulations of large systems such as proteins and polynucleotides, MM force fields are well established and widely used. Established MM force fields rely on lookup tables with a finite set of atom types characterized by chemical properties of the atom and its bonded neighbors for parameterization. Recently, the Espaloma approach (Wang et al., 2022) has demonstrated that machine learning can be used to increase the accuracy of MM force fields by learning to assign parameters based on a graph representation of the molecule with chemical properties that rely on expert knowledge, such as orbital hybridization states or formal charge, as input features.
@@
In this work, we propose a novel machine learning framework, Grappa (Graph Attentional Protein Parametrization), to learn MM parameters directly from the molecular graph, improving accuracy on a broad range of chemical space and eliminating the need for hand-crafted features. Grappa employs a graph attentional neural network to construct atom embeddings capable of representing chemical environments based on the 2D molecular graph, followed by a transformer (Vaswani et al., 2023) with symmetry-preserving positional encoding. Since MM parameters only have to be predicted once per molecule, Grappa can be incorporated into highly optimized MM engines such as GROMACS (Abraham et al., 2015) and OpenMM (Eastman et al., 2013). This allows energy and force evaluations with the same computational cost as traditional force fields, at state-of-the-art MM accuracy.
@@
We show that Grappa outperforms traditional MM force fields and the machine-learned MM force field Espaloma on the Espaloma dataset (Takaba et al., 2023), which contains over 14,000 molecules and more than one million states, covering small molecules, peptides and RNA. Since Grappa uses no hand-crafted chemical features, it can be extended to uncharted regions of chemical space, which we demonstrate on the example of peptide radicals. Grappa is transferable to individual macromolecules and assemblies such as proteins and viruses, which exhibit similar dynamics as established force fields. Starting from an unfolded initial state, MD simulations of small proteins parametrized by Grappa recover experimentally determined folding structures of small proteins, suggesting that Grappa captures the physics underlying protein folding. We demonstrate the efficiency of Grappa, which is inhereted from MM, by simulating a system of one million atoms with the proposed force field on a single GPU, with a similar number of timesteps per second as a highly performant E(3) equivariant neural network (Musaelian et al., 2022) on over 4,000 GPUs.
@@ 1.1 Molecular Mechanics
In MM, the potential energy of a system with a given molecular graph is expressed as a sum of contributions from different interactions. Bonded interactions are described by functions of E(3)-invariant internal coordinates such as the lengths $r_{ij}$ of bonds between two atoms, angles $\theta_{ijk}$ between three consecutive atoms and dihedrals $\phi_{ijkl}$ of two planes spanned by four atoms. For the dihedrals, one considers interactions between four atoms that are either consecutively bonded (torsions) or where three atoms are bonded to a central atom (impropers), which do not reflect an independent degree of freedom but are used to maintain planarity of certain chemical groups. One commonly uses harmonic potentials for bond stretching and angle bending and a periodic function for the dihedral potential. The potential energy of all interactions along bonds then is given by
@@
For the periodic dihedral potential, a common choice is a Fourier series with the constraint that the dihedral potential is extremal at and symmetric around zero, which eliminates the need for sine terms. Additionally, atom pairs that are not included in such N-body bonded interaction terms contribute to the potential energy through pairwise nonbonded interaction, typically described by Lennard-Jones and Coulomb potentials.
@@ 2 Grappa
Inspired by the atom typing with hand-crafted rules in traditional MM force fields and in analogy to Wang et al. 2022, Grappa first predicts $d$-dimensional atom embeddings,
@@
Since the mapping from molecular graph to energy is differentiable with respect to the model parameters and spatial positions, it can be optimized on predicting QM energies and forces end-to-end, as visualized in Figure 3. Notably, the machine learning model does not depend on the spatial conformation of the molecule, thus it has to be evaluated only once per molecule and the computational cost of each subsequent energy evaluation is given by the MM energy functional.
Grappa currently only predicts bonded MM parameters since we expect that nonbonded interactions are not covered sufficiently by the monomeric datasets used for training, rendering the nonbonded parameters underdetermined. The nonbonded parameters are taken from established MM force fields that can reproduce solute interactions and melting points, which we expect to be strongly dependent on nonbonded interactions.
@@ 2.1 Permutation symmetries in MM
For the mapping from atom embeddings $\bm{\nu}$ to MM parameters, we postulate certain permutation symmetries that the model should respect. To derive these symmetries, we consider the energy function of MM as a decomposition into contributions from subgraphs of the featurized molecular graph that correspond to bonds, angles, torsions and improper dihedrals. We demand invariance of the energy contribution under node permutations that induce isomorphisms of the respective subgraph. For bonds, angles and torsions, these permutations leave the respective spatial coordinate invariant, thus we can achieve invariance of the energy contribution by demanding invariance of the MM parameters,
@@ 2.2 The Grappa architecture
To predict atom embeddings from the molecular graph, Grappa employs a graph attentional neural network inspired by the transformer architecture (Vaswani et al., 2023). Multi-head dot-product attention on graph edges (Veličković et al., 2018) is followed by feed-forward layers with residual connections (He et al., 2015) and layer normalization (Ba et al., 2016), which has been demonstrated to enhance the expressivity of attention layers (Brody et al., 2023).
@@
defining a symmetry pooling operation with $\mathcal{P}$-invariant output. We call this combination of permutation invariant positional encoding with permutation equivariant layers and symmetric pooling the symmetric transformer, which is illustrated in Figure 2. The symmetric transformer can be generalized to permutation symmetries of arbitrary subgraphs by using the eigenvectors of the graph Laplacian (Yun et al., 2020) as positional encoding.
Finally, we map to the range of physically sensible parameters, e.g. $(0,\infty)$ for bond and angle force constants or $(0,\pi)$ for equilibrium angles $\theta^{(0)}$. To this end, we use scaled and shifted versions of ELU and the sigmoid function as described in A.2.3. While one could also use the exponential for mapping to $(0,\infty)$, ELU’s linear behaviour towards large inputs is favorable for producing stable gradients during optimization. With the scaling we ensure that a normally distributed output of the neural network is mapped to a distribution with mean and standard deviation that is suitable for the respective MM parameter, which can be seen as a normalization technique (Ba et al., 2016). For predicting dihedral parameters $\xi_{ijkl}$, we use a sigmoid gate, which allows the model to supress dihedral modes that are not needed.
As input feature, we use one-hot encodings of the atomic number, the number of neighbors of the respective node and membership in loops of length 3 to 8, which can be directly calculated from the molecular graph. The assignment of nonbonded MM parameters is done using a traditional force field of choice. Since the bonded parameters predicted by Grappa may depend on the scheme by which nonbonded parameters are assigned, we pass the partial charge of each atom as input feature to Grappa, which also allows to encode the total charge of a molecule without breaking graph symmetries as one would potentially do by using the formal charge instead.
@@ 2.3 Training
Molecular GraphPositionsQM EnergyQM ForceGrappaMM ParametersMM EnergyFunctionalPred. EnergyPred. ForceLossPositionsPred. EnergyPred. ForceQM EnergyQM ForceBackprop.Inference
Figure 3: Grappa predicts one set of parameters per molecule. With the MM energy functional (Eq. 5), the parameters can be mapped to energies and forces of given states, whose deviation from the ground truth is minimized during training. State-specific quantities are represented in green, molecule-specific quantities are represented in grey.
We train Grappa to minimize the mean squared error (MSE) between QM energies $E_{\text{QM}}$ and forces $-\nabla_{\mathbf{x}}E_{\text{QM}}$, whose contribution we weight by the hyperparameter $\lambda_{F}$, and its prediction, $E$ and $\nabla_{\mathbf{x}}E$ respectively, which includes the non-learnable nonbonded contribution. At the start of training, we also include the deviation of predicted bond, angle and torsion MM parameters to those of a given traditional force field, weighted by $\lambda_{\text{trad}}$ and, for regularization, the $L2$ norm of dihedral MM parameters $\xi_{ijkl}^{(\text{dih})}$, weighted by $\lambda_{\text{dih}}$ as in (Takaba et al., 2023). That is, we use the loss function
@@
Since MM can only predict energy differences of states, not formation energies, we subtract the mean of target and predicted energies for each molecule. As shown previously for other machine learned force fields (Schütt et al., 2017, 2021), we find training on forces $-\nabla_{r}E$ in addition to energies important also in our setting, for two reasons. First, the energy is a global, pooled quantity and thus less expressive than the forces, which are local and by a factor of $3N$ more numerous. Second, as shown in (Czarnecki et al., 2017), learning derivatives of the target with respect to the input can lead to improved generalization and data efficiency, effectively smoothing the learned potential energy surface.
Details on the training procedure and hyperparameters can be found in A.3.
@@ 3 Results
3.1 Grappa is state-of-the-art
To demonstrate that Grappa is the current state-of-the-art MM force field in terms of accuracy, we train and evaluate a Grappa model on the dataset reported in the follow-up paper (Takaba et al., 2023) to Espaloma, which contains 17,427 unique molecules and over one million conformations. Our training and test partition is identical with the one from Espaloma, where the molecules were divided into 80% training, 10% validation and 10% test set, based on isomeric SMILES strings. The dataset covers small molecules, peptides and RNA with states sampled from the Boltzmann distribution at 300 K and 500 K, from optimization trajectories and from torsion scans. The nonbonded contribution is calculated using the OpenFF-2.0.0 force field (Boothroyd et al., 2022) and partial charges from the AM1-BCC method, as in Espaloma. We train the model for 1,000 epochs on an A100 GPU, which takes about one day.
For all types of molecules, Grappa outperforms established MM force fields and Espaloma in terms of energy and force accuracy on Boltzmann-sampled states as shown in Table 1 and Figure 4a. While the Boltzmann samples are the more relevant benchmarking data for using Grappa in MD simulations, we also compare performance on torsion scans and optimization trajectories (Table A6). There, we find Grappa and Espaloma to be competitive, while Grappa is more accurate for forces and less accurate for energies. Grappa also outperforms the baselines if only trained on a fraction of Espaloma’s training set (Figure A10), indicating high data efficiency. Chemical properties based on expert knowledge, e.g. hybridization and aromaticity, have long been used to assign MM parameters (He et al., 2020; Boothroyd et al., 2022). We show that accurate MM parameters can be predicted directly from the molecular graph, without relying on hand-crafted input features.
@@ 3.1 Grappa is state-of-the-art
Table 1: Accuracy of Grappa, Espaloma 0.3 (Takaba et al., 2023) and established MM force fields on test molecules of the Espaloma dataset. We report the RMSE of centered energies in kcal/mol and the componentwise RMSE of forces in kcal/mol/Å, uncertainties can be found in Table A6. Gaff-2.11 (He et al., 2020) is a general-purpose force field, ff14SB (Maier et al., 2015) is an established protein force field and RNA.OL3 (Zgarbová et al., 2011) is specialized to RNA. The SPICE-Pubchem and SPICE-DES-Monomers datasets, containing small molecules, along with the dipeptide dataset, are subsets of the SPICE dataset (Eastman et al., 2023). The RNA datasets feature trinucleotide states calculated with the B3LYP-D3BJ functional.
@@
Figure 4: Comparison of energy predictions of (a) Grappa-1.3 and the established force fields (b) Gaff-2.11, ff99SB-ILDN and RNA.OL3 for test molecules from Espaloma’s SPICE-Pubchem, SPICE-Dipeptide and RNA-Trinucleotide datasets; force predictions are depicted at A12. (c) The first principal components $u_{1}$ and $u_{2}$ of predicted atom embeddings from the Espaloma test dataset can be related to a combination of the main group and period in the periodic table of elements. Lines of constant main group or period are represented by approximate diagonals in latent space.
3.2 Grappa is extensible across chemical space